OpenStudy (anonymous):

Find the average rate of change for f(x) = √x+3 from 1 to 2 and the equation of the secant line containing (1, f(1)) and (2, f(2)). Please show all work. Thanks! Find the average rate of change for f(x) = √x+3 from 1 to 2 and the equation of the secant line containing (1, f(1)) and (2, f(2)). Please show all work. Thanks! @Mathematics

OpenStudy (anonymous):

rate of change (which is also the slope)= change in f(x)/change in x =(f(2)-f(1))/(2-1)

OpenStudy (anonymous):

no no no

OpenStudy (anonymous):

this is wrong

OpenStudy (anonymous):

r bliss is wrong

OpenStudy (anonymous):

well Iwasnnot finished. what do you think is wrong with what I did so are

OpenStudy (anonymous):

gotta divide the divedt by 7 ya herd, then u mulitply the coherent reply, #FRIED

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

since you have a little trouble typing I have to say that I have no idea what you are trying to say.

OpenStudy (anonymous):

its fairly simple, acording to several goeologists it can be co hesive and rather peprotory. Therefroe in order to find the sum of the equation you must use several varibales of defense. Eqaully submissive but nerologgically ttrasmitted.

OpenStudy (anonymous):

sleep it off.

OpenStudy (anonymous):

u were proved wrong :p

OpenStudy (anonymous):

swag

OpenStudy (anonymous):

Hey can you guys provide me a solution to this rather than argue here? lol Thank you.

OpenStudy (anonymous):

loooooooooooooooooooool

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

iight im bck

OpenStudy (anonymous):

so basically heres the solution

OpenStudy (anonymous):

...

OpenStudy (anonymous):

basically this is a paraphyletic group of organisms that consist of all gill-bearing aquatic vertebrate (or craniate) animals that lacks limbs with digits. Included in this definition are the living hagfish, lampreys, and cartilaginous and bony fish, as well as various extinct related groups. Most fish are "cold-blooded", or ectothermic, allowing their body temperatures to vary as ambient temperatures change, though some of the large active swimmers like white shark and tuna can hold a higher core temperature.[1][2] Fish are abundant in most bodies of water. They can be found in nearly all aquatic environments, from high mountain streams (e.g., char and gudgeon) to the abyssal and even hadal depths of the deepest oceans (e.g., gulpers and anglerfish). At 32,000 species, fish exhibit greater species diversity than any other class of vertebrates.[3] Fish, especially as food, are an important resource worldwide. Commercial and subsistence fishers hunt fish in wild fisheries (see fishing) or farm them in ponds or in cages in the ocean (see aquaculture). They are also caught by recreational fishers, kept as pets, raised by fishkeepers, and exhibited in public aquaria. Fish have had a role in culture through the ages, serving as deities, religious symbols, and as the subjects of art, books and movies. Because the term "fish" is defined negatively, and excludes the tetrapods (i.e., the amphibians, reptiles, birds and mammals) which descend from within the same ancestry, it is paraphyletic, and is not considered a proper grouping in systematic biology. The traditional term pisces (also ichthyes) is considered a typological, but not a phylogenetic classification.

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

hold on, Im doing math- not science....

OpenStudy (anonymous):

rate of change (which is also the slope)= change in f(x)/change in x =(f(2)-f(1))/(2-1) is the function the (square root of x) +3 or the squre root of (x+3)? I think it must be the squre root of (x+3). so =(sq root of 5 -sq root of 4)/1 This is the slope. the equation of the secan t line is f(x) = slope *x + constant constant = f(x) - slope*x for x=1 =2-(sq rt (5)-2)/1*1 =4-sq rt(5)

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

well did I guess the function that you meant correctly? I can't tell from what you wrote where the sq root sign ends.

OpenStudy (anonymous):

sqrt sign ends over the 3. sorry

OpenStudy (anonymous):

\[f(x)= \sqrt{x+3}\] this is what it should be